OpenStudy (anonymous):

-x^3+3x+1=0 @phi can you help me with this too please

OpenStudy (solomonzelman):

There is a formula for the exact answer, and it looks very weird. I don't remember it.

OpenStudy (phi):

are you sure you have a cubic ? they make quadratics look trivial

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i used a calculator on google and the answer looked really weird had inside a sqrt fractions multiplying

OpenStudy (phi):

this is a "depressed cubic" because it does not have an x^2 term. We can try https://en.wikipedia.org/wiki/Cubic_equation#Vieta.27s_substitution

OpenStudy (anonymous):

where did the w come from in that equation?

OpenStudy (phi):

you should multiply through by -1 to get x^3 -3x - 1 = 0 and match to x^3 + p x + q=0 to find p= -3 and q= -1 now we solve w^6 + q w^3 - p^3/27 =0 with our numbers w^6 - 3 w^3 +1/27 = 0 let u= w^3, and solve u^2 -3u + 1/27 = 0

OpenStudy (anonymous):

would i simplify the 27?

OpenStudy (anonymous):

\[3^{3}\]

OpenStudy (phi):

wait. I got my p's and q's mixed up (I did not "mind by p's and q's"

OpenStudy (phi):

now we solve w^6 + q w^3 - p^3/27 =0 p= -3 and q= -1 so we should be doing w^6 - w^3 - (-3)^3/27 = 0 w^6 - w^3 + 1 = 0

OpenStudy (phi):

let u= w^3 u^2 - u + 1 = 0

OpenStudy (anonymous):

ok im behind a bit where did the P,Q,and W comes from

OpenStudy (phi):

I am (trying to ) doing this https://en.wikipedia.org/wiki/Cubic_equation#Vieta.27s_substitution

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

\[u ^{2}-u=-1\]

OpenStudy (anonymous):

do i divide the -u?

OpenStudy (anonymous):

\[u ^{2}=\frac{ 1 }{ u }\]

OpenStudy (anonymous):

then sqrt ?\[u=\sqrt{\frac{ 1 }{ u }}\]

OpenStudy (phi):

either do the quadratic formula or complete the square.

OpenStudy (anonymous):

the Quadratic formula is \[x=\frac{ -b \sqrt{b ^{2}-4ac} }{ 2a }\] correct with A=1 B=-1 and c=1?

OpenStudy (anonymous):

\[x=\frac{ 1\sqrt{-1^{2}-4(1)(1)} }{ 2(1) }\]

OpenStudy (anonymous):

\[x=\frac{ 1\sqrt{-3} }{ 2 }\]

OpenStudy (anonymous):

\[x=\pm \frac{ -1\sqrt{-3} }{ 2 }\]

OpenStudy (anonymous):

oops \[x=\pm \sqrt{\frac{ -1\sqrt{-3} }{ 2 }}\]

OpenStudy (phi):

you got confused. it is (-b ± sqr(-3) )/2 or (1 ± sqr(-3))/2

OpenStudy (anonymous):

so \[x=\frac{1\pm \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

didnt think the other one looked right

OpenStudy (phi):

u^2 - u + 1 = 0 a=1, b= -1, c=1 yes \[ x=\frac{1\pm \sqrt{3} }{ 2 } \] we now need the cube root of this ( gack!)

OpenStudy (phi):

oh, the sqr(-3) is sqr(3) * i where i = sqr(-1)

OpenStudy (phi):

we got a complex number. we may be wandering into stuff you don't know? in polar form, \[ x = e^{i \frac{\pi}{3}} \] the cube root is \[ x^{\frac{1}{3}} = e^{i \frac{\pi}{3}\cdot \frac{1}{3}} =e^{i \frac{\pi}{9}} \]

OpenStudy (phi):

according to https://en.wikipedia.org/wiki/Cubic_equation#Vieta.27s_substitution one root is t1= x - p/(3x) p is -3 and x is e^(i pi / 9) and we get t1= e^(i pi/9) + 3/3e^(i pi / 9) = e^(i pi/9) + e^(-i pi/9) = 2 cos( pi/9)

OpenStudy (anonymous):

-0.2repeating

OpenStudy (phi):

? one root is 2 cos(20º) = 1.87938524 the other two roots are equally spaced 120º apart 2 cos(120+20) = 2 cos(140)= 1.5321 2 cos(240+20) = 2 cos(260) = -0.3473

OpenStudy (anonymous):

oh i just typed 2 cos pi over 9 into google

OpenStudy (phi):

oh. you should put in parens 2*cos(pi/9) because it did 2*cos(pi) / 9 which is different

OpenStudy (anonymous):

1.87938524157

OpenStudy (phi):

in radians the other two roots are 2 cos( 2pi/3 + pi/9) = 2 cos(7 pi/9)= -1.5321 (I left a minus sign out up above ) and 2 cos (4pi/3 + pi/9) = 2 cos(13 pi/9) = -0.3472963

OpenStudy (anonymous):

im confused now whats all this?

OpenStudy (phi):

the cubic has 3 solutions

OpenStudy (anonymous):

ohh i forgot about that

OpenStudy (phi):

according to https://en.wikipedia.org/wiki/Cubic_equation#Vieta.27s_substitution the roots are t= x - p/(3x) with p= -3 we get t= x + 1/x or t= x + x^(-1) when x is exp(i theta) this turns into 2 cos(theta) t= 2 cos(theta) where x is the cube root of \[ u = \frac{1\pm \sqrt{3}i }{ 2 } \] I found one cube root of u as e^( i*pi/9) the other two cube roots are separated by 120º (2 pi/ 3 radians) in other words we have x= exp( i pi/9) x= exp( i 7pi/9) x= exp(i 13 pi/9)

OpenStudy (phi):

these seems awfully complicated stuff.

OpenStudy (anonymous):

very complicated never done anything like this